OpenStudy (anonymous):

During the first part of a trip, a canoeist travels 57 miles at a certain speed. The canoeist travels 18 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 5 hours. What was the speed(mph) on each part of the trip?

OpenStudy (anonymous):

\[velocity=\frac{distance}{time}\] First we can find the average v for the whole travel. v=(57+18)/5=75/5=15 mph How can we use that ?

OpenStudy (anonymous):

your guess is as good as mines i have no clue and just need help on the answer i have been stuck on this problem for 30 min.

OpenStudy (lgbasallote):

the canoist travels 57 miles at a certain speed let n be this speed the time he traveled here is 57/n (because time = distance/speed) the canoeist travels 18 miles at a speed 5 mph slower than befor let n - 5 be his speed at this interval the time he traveled here is 18/n-5 the total time is 5 hours therefore \[\frac{57}{n} + \frac{18}{n-5} = 5\] does that make sense?

OpenStudy (anonymous):

lol this is what i got using a similar question i asked earlier but i need to know now how to convert this into a decimal to determine the mph i got 3.44256 and 16.5574

OpenStudy (lgbasallote):

determine what into decimal?